OpenStudy (anonymous):

a merchant wants to stock the shelves with 100 Ibs. of mixed nuts that cost $8.00 a Ibs. the merchant has cashews and peanuts. peanuts cost $6.00 a Ibs. and cashews that cost $12.00 a Ibs. how many Ibs. of each should the merchant use in the mixture?

OpenStudy (anonymous):

x+y=100 lbs x =100-y X for cashews y for peanuts ($12*X)+($6*y)=(100*80) Subtitution X =100-Y ($12*(100-y))+($6*y)=$800 1200-12y +6y =800 -6y =800-1200 y =-400/-6 y = 66.7 X+Y=100bls x+ 66.7 =100 X=100-66.7 X=33.3 you may round : 33.3= 33 , 66.7 =67 33 bls cashew 67 bls peanut

OpenStudy (anonymous):

thank you soooo much